OpenStudy (anonymous):

find the probabilities for each, using the standard normal distrubition.. * P(1.12

OpenStudy (anonymous):

0.055 5.5%

OpenStudy (anonymous):

\[\int_{1.12}^{1.43} \frac{1}{\sqrt{2\pi}}e^{-x^2/2} dx \approx 0.55.\]